# In Search of a Hidden Curve

### v1

@@ Acknowledgement.
The author would like to express his hearty gratitude to the organizer, Laura P. Schaposnik, of the Workshop and School on Complex Lagrangians, Integrable Systems, and Quantization, which was held at the Mathematical Institute of the University of Oxford in summer 2023, for her invitation and kind hospitality, and her continued support and encouragement to him over the past decade. His thanks are due to her co-organizers Mengxue Yang and Christopher Mahadeo as well for their effort to make the workshop successful. He thanks Zachary Ibarra for many discussions on the topics of $\zeta(3)$, which were very useful to compile Section 4 of these notes.
@@ 1.1. Theme $1$: Mirror symmetry of Catalan numbers
The first aim of these lectures is to present a few problems that have been solved when a hidden curve behind the scenes was discovered. These curves are commonly called spectral curves. As we see below, as soon as the spectral curve is identified, it acts as a catalyst to let the subject in question weave the whole story in front of us. And the important common feature of these curves is that they are complex Lagrangians in a complex symplectic surface.
@@
(1.1) $\Sigma=\left\{(x,z)\;\left|\;x=z+\frac{1}{z}\right.\right\}$
@@
$C_{m}=\frac{2(2m-1)}{m+1}C_{m-1},\qquad C_{0}=1,$
with respect to the generating function $z(x)$. The differential equation (1.3) has three regular singular points at $x=-2,2,\infty$. (The definition of this terminology is given below in Definition 1.2.) The analytic continuations of $z(x)$ to ${\mathbb{C}}$ from the neighborhood of $\infty\in{\mathbb{P}}^{1}$ are
@@
Even though this example is trivial in many sense, the point here is that (1.3) is a Picard-Fuchs equation associated with the projection $\pi:\Sigma\longrightarrow{\mathbb{P}}^{1}$ defined by the algebraic equation $x=z+1/z$. This is a trivial Landau-Ginzburg model. A coordinate transformation $x=4t-2$ brings this differential equation to one of the Euler-Gauß hypergeometric differential equations,
@@
be a second order differential equation defined around a neighborhood of $x=0$ on a small disc $|x|<\epsilon$ with meromorphic coefficients $a_{1}(x)$ and $a_{2}(x)$ with poles at $x=0$. Denote by $k$ (reps. $\ell$) the order of the pole of $a_{1}(x)$ (resp. $a_{2}(x)$) at $x=0$. If $k\leq 1$ and $\ell\leq 2$, then (1.6) has a regular singular point at $x=0$. Otherwise, consider the Newton polygon of the order of poles of the coefficients of (1.6). It is the upper part of the convex hull of three points $(0,0),(1,k),(2,\ell)$. As a convention, if $a_{j}(x)$ is identically 00, then we assign $\infty$- ∞ as its pole order. Let $(1,r)$ be the intersection point of the Newton polygon and the line $x=1$, $r=\begin{cases}k\qquad 2k\geq\ell,\\ \frac{\ell}{2}\qquad 2k\leq\ell.\end{cases}$ If $r>1$, (1.6) has an irregular singular point of class $r-1$.
@@
Figure 1.1. $1$-Skeleton of a Cell Decomposition and a Cell Graph
Here comes the relation to enumeration. The coefficients of the expansion $C_{g,n}(\mu_{1},\dots,\mu_{n})$ of (1.10) are the generalized Catalan numbers of type $(g,n)$ that count the numbers of cell graphs of genus $g$ and $n$ labeled vertices of degrees $(\mu_{1},\dots,\mu_{n})$ [30, 31, 33, 71, 72]. A cell graph is the $1$-skeleton of a cell decomposition of a connected oriented surface of genus $g$ with $n$ labeled 00-cells. Its dual graph is commonly called a ribbon graph on which faces are labeled. To avoid the difficulties of counting automorphisms of graphs, we impose that no cyclic rotations of incident edges at any vertex are allowed, as in Figure 0.2. This explains the denominator $\mu_{1}\cdots\mu_{n}$ of (1.10) (see Remark 2.29). For $(g,n)=(0,1)$, the unique vertex has to have an even degree, and $C_{0,1}(2m)=C_{m}$ is the $m$-th Catalan number. The surprising properties of the functions $F_{g,n}$ are the following (see for example, [31, Theorem 2.7]).
@@
For the case of $02g=2+n>$ 0, substitute each $x_{i}$ in (1.10) with $\begin{cases}x_{i}=z_{i}+\frac{1}{z_{i}},\\ z_{i}=\frac{t_{i}+1}{t_{i}-1},\end{cases}$ and write the result as $F_{g,n}(t_{1},\dots,t_{n})$ incorporating these substitutions. Then for every $(g,n)$ in the range of $02g-2+n>$ 0, the following holds:
  • $F_{g,n}(t_{1},\dots,t_{n})$ is a Laurent polynomial in the $t$-variables of degree $6g-6+3n$.
@@
Here, ${\mathcal{M}}_{g,n}$ is the moduli space of smooth $n$-pointed algebraic curves of genus $g$, ¯${\overline{\mathcal{M}}}_{g,n}$ its compactification, i.e., the Deligne-Mumford stack of stable curves of finite type $(g,n)$, and
@@
Although we do not touch in these lectures, there is yet another coordinate change from $(t_{1},t_{2},t_{3},\dots)$ appearing in the Catalan case to the time evolution parameters $(t_{1},t_{2},t_{3},\dots)$ in the Lax equations (1.17) defined below. Our use of the same notations is due to the conventions of the author’s previous publications. In reality, these are not the same variables. After the choice of the right coordinate transformation,
@@ 1.2. Theme $2$: When is a space a moduli space?
A moduli space is different. On this space, every point has its unique name. Each point is distinguished by its individuality. No points on a moduli space are the same. A neighborhood of each point is different from place to place. There may also be singularities on a moduli space. Each point has its own characteristic importance, yet it is a member of a smooth family. Since every point is different, the moduli space itself presents its miracle: it can exist, as a harmonious space!
@@
We give a linear coordinate system on ${\mathbb{C}}^{n}$ and represent the $2n$ generators of $\Gamma$ by a set of $2n$ column vectors $w_{1},\dots,w_{2n}$ of size $n$. They form a complex matrix $\Pi$ of size $n\times 2n$. $GL_{n}({\mathbb{C}})$ acts on $\Pi$ from the left, representing the coordinate change. Since the column vectors of the matrix $\Pi$ generate the ${\mathbb{Z}}$-module $\Gamma$, $SL_{2n}({\mathbb{Z}})$ acts on $\Pi$ from the right, representing change of generators of the same lattice $\Gamma$. Since $\{w_{1},\dots,w_{2n}\}$ is linearly independent over ${\mathbb{R}}$, there is a ${\mathbb{C}}$-linearly independent subset consisting of $n$ vectors. Thus we can use a left multiplication of $GL_{n}({\mathbb{C}})$ and a right multiplication of $SL_{2n}({\mathbb{Z}})$ on $\Pi$ to bring it to the special shape $[I\,|\,\Omega]$. The geometry of the quotient space ${\mathbb{C}}^{n}/\Gamma$ is encoded in this $n\times n$ matrix $\Omega$, which we call the period matrix of the torus $T={\mathbb{C}}^{n}/\Gamma$.
@@
is the moduli space of holomorphic line bundles on $C$ of degree 00. For each fixed $g>1$, there are $33g-$ 3 dimensional family of Jacobian varieties, while Abelian varieties form a family of $g(g+1)/2$ dimensions. Therefore, Jacobians are very special among Abelian varieties.
@@ 1.3. Theme $3$: Finding a curve from a period matrix
(1.16) $\vartheta(z,\Omega)=\sum_{n\in{\mathbb{Z}}^{n}}\exp\left(2\pi i{\langle}n,z{\rangle}+\pi i{\langle}n,\Omega n{\rangle}\right),\qquad z\in{\mathbb{C}}^{n},$
[…]
@@
(1.17) $\frac{\partial L}{\partial t_{n}}=[B_{n},L],\qquad B_{n}=(L^{n})_{+},\quad n=1,2,3,\dots,$
@@
(1.18) $L=\partial+\sum_{i=1}^{\infty}u_{i+1}(x;\mathbf{t})\partial^{-i},\qquad\partial=\frac{\partial}{\partial x},\quad\mathbf{t}=(t_{1},t_{2},t_{3},\dots)$
@@
which does not depend on $y$ gives a solution to the KP equation. A relation between the KdV equation (1.19) and algebraic curves is easily seen by considering a traveling wave solution with speed $-c$ constructed from the Weierstraß  elliptic function $\wp(z)$:
@@ 1.4. Theme $4$: The $\tau$-functions
on the Sato Grassmannian, where $V={\mathbb{C}}((z))$ and $\Lambda\in\mathfrak{gl}_{\infty}$ is the maximal nilpotent element corresponding to the multiplication of $z^{-1}$ on $V$ (see for example, [70, Chapter 7]). The $\tau$-function records this action through the vantage point of the determinant line bundle. This also explains why so many soliton equations are exactly solved in terms of determinant. Since the initial value problem of the KP equations is uniquely solvable ([68, 69]), the space of solutions to the KP equations is identified with $Gr(\mu=0)$.
@@
the space of meromorphic sections of ${\mathcal{L}}$ with arbitrary poles at $p$, where $p\in C$ is a non-singular point of $C$, $z$ is a local parameter of $C$ around $p$, and ${\mathcal{L}}$ is a line bundle on $C$ of degree $g(C)-1$. We then have (see [68, Lemma 3.7])
@@
2. Fugue of the theme of spectral curves
@@ 2.1. The Laplace transform
This conversation took place in the lobby of MSRI, Berkeley, after the end of the day. We shook hands tightly and happily, as we started to depart. We then noticed that there was a young mathematician listening our conversation. At the end, he shouted:
@@
A mathematical picture has emerged in the last two decades since the discoveries of Eynard-Orantin [37], Mariño [65], Bouchard-Klemm-Mariño-Pasquetti [13] and Bouchard-Mariño [14] in physics, and many mathematical efforts including [12, 17, 33, 34, 36, 40, 61, 72, 73, 75, 76, 79, 80, 81, 82, 83, 84, 90]. As a working hypothesis, we phrase it in the form of a principle.
@@
There are many examples of mathematical problems that fall in to this principle. Among them is the theory of simple Hurwitz numbers [36, 76] that we are going to present now. Besides Hurwitz numbers, numerous mathematical results have been established. They include counting of Grothendieck’s dessins d’enfants [17, 72, 79, 80, 83], higher genus Catalan numbers [31, 33], orbifold Hurwitz numbers [12], double Hurwitz numbers and higher spin structures [74], and the stationary Gromov-Witten invariants of ${\mathbb{P}}^{1}$ [34, 84]. One of the most important results in the Gromov-Witten setting is the remodeling conjecture [13] and its solution [40] mentioned above. A new direction was suggested in Kontsevich-Soibelman [56]. It is now impossible to list all research in this direction, simply called topological recursion, so we refer to recent papers by many authors and the papers they cite: Andersen, Borot, Bouchard, Chidambaram, Do, Dunin-Barkowski, Eynard, Garcia-Failde, Iwaki, Lewański, Norbury, Orantin, Osuga, Shadrin, … A comprehensive and the most recent introduction is given by Bouchard [11], in which one can see how the focus of the topological recursion community has changed into new developments in the last decade.
@@
2.2. Simple Hurwitz numbers [51]
In this subsection, we define simple Hurwitz numbers, and a combinatorial equation that they satisfy, the cut-and-join equation, is proved. In the next subsection, the Laplace transform of the Hurwitz numbers is presented. The Laplace transformed holomorphic functions live on the mirror $B$-side of the model, according to Princeple 2.3. The Lambert curve is defined as the domain of holomorphy of these holomorphic functions. Then in the following subsections, we give the Laplace transform of the cut-and-join equation. The result is a simple polynomial recursion formula and is equivalent to the Eynard-Orantin topological recursion for the Lambert curve. We also give a straightforward and simple derivation [76] of the Witten-Kontsevich theorem on the $\psi$-class intersection numbers [23, 54, 95], and the $\lambda_{g}$-formula of Faber and Pandharipande [38, 39], using the recursion formula we establish in [76].

### v2

@@ Acknowledgement.
The author would like to express his hearty gratitude to the organizer, Laura P. Schaposnik, of the Workshop and School on Complex Lagrangians, Integrable Systems, and Quantization, which was held at the Mathematical Institute of the University of Oxford in summer 2023, for her invitation and kind hospitality, and her continued support and encouragement to him over the past decade. His thanks are due to Schaposnik’s co-organizer and co-editor of the volume Mengxue Yang for her effort to make the workshop and the volume successful. He thanks Zachary Ibarra for many discussions on the topics of $\zeta(3)$, which were very useful to compile Section 4 of these notes.
@@ 1.1. Theme $1$: Mirror symmetry of Catalan numbers
The first aim of these lectures is to present a few problems that have been solved when a hidden curve behind the scenes is identified. These curves are commonly called spectral curves. As we see below, as soon as the spectral curve is identified, it acts as a catalyst to let the subject in question weave the whole story in front of us. And the important common feature of these curves is that they are complex Lagrangians in a complex symplectic surface.
Spectral curves have appeared in many different contexts in the past, including integrable systems of nonlinear PDEs such as the KdV equations, Hitchin’s work on Higgs bundles, and analysis of random matrices. In each situation, there is a concrete definition of what people call a spectral curve. In these lectures, however, we do not attempt to give a universally valid definition of spectral curves, except that we say, a spectral curve is a collection of eigenvalues of an operator that takes the shape of a curve. The fact that we have a spectral curve for our counting problem suggests that it has a hidden connection to integrable systems, algebraic geometry, random matrix theory, and beyond.
@@
[…]
@@
$C_{m}=\frac{2(2m-1)}{m+1}C_{m-1},\hskip 23.49976ptC_{0}=1,$
with respect to the generating function $z(x)$. The differential equation (1.3) has three regular singular points at $x=-2,2,\infty$. (The definition of this terminology is given below in Definition 1.2.) The analytic continuations of $z(x)$ from the neighborhood of $\infty\in{\mathbb{P}}^{1}$ are
@@
Even though this example is trivial in many sense, the point here is that (1.3) is a Picard-Fuchs equation associated with the projection $\pi:\Sigma\longrightarrow{\mathbb{P}}^{1}$ defined by the algebraic equation $x=z+1/z$. This is a trivial Landau-Ginzburg model. A coordinate transformation x=4t2x=4t-2 brings this differential equation to one of the Euler-Gauß hypergeometric differential equations,
@@
be a second order differential equation defined around a neighborhood of $x=0$ on a small disc $|x|<\epsilon$ with meromorphic coefficients $a_{1}(x)$ and $a_{2}(x)$ with poles at $x=0$. Denote by $k$ (reps. $\ell$) the order of the pole of $a_{1}(x)$ (resp. $a_{2}(x)$) at $x=0$. If $k\leq 1$ and $\ell\leq 2$, then (1.6) has a regular singular point at $x=0$. Otherwise, consider the Newton polygon of the order of poles of the coefficients of (1.6). It is the upper part of the convex hull of three points $(0,0),(1,k),(2,\ell)$. As a convention, if $a_{j}(x)$ is identically 0, then we assign -$\infty$ as its pole order. Let $(1,r)$ be the intersection point of the Newton polygon and the line $x=1$, $r=\begin{cases}k\qquad 2k\geq\ell,\\ \frac{\ell}{2}\qquad 2k\leq\ell.\end{cases}$ If $r>1$, (1.6) has an irregular singular point of class r1r-1.
@@
Figure $1.$ 11-Skeleton of a Cell Decomposition and a Cell Graph
Here comes the relation to enumeration. The coefficients of the expansion $C_{g,n}(\mu_{1},\dots,\mu_{n})$ of (1.10) are the generalized Catalan numbers of type $(g,n)$ that count the numbers of cell graphs of genus $g$ and $n$ labeled vertices of degrees $(\mu_{1},\dots,\mu_{n})$ [31, 32, 34, 75, 76]. A cell graph is the $1$-skeleton of a cell decomposition of a connected oriented surface of genus $g$ with $n$ labeled 0-cells. Its dual graph is commonly called a ribbon graph on which faces are labeled. To avoid the difficulties of counting automorphisms of graphs, we impose that no cyclic rotations of incident edges at any vertex are allowed, as in Figure 0.2. This explains the denominator $\mu_{1}\cdots\mu_{n}$ of (1.10) (see Remark 2.29). For $(g,n)=(0,1)$, the unique vertex has to have an even degree, and $C_{0,1}(2m)=C_{m}$ is the $m$-th Catalan number. The surprising properties of the functions $F_{g,n}$ are the following (see for example, [32, Theorem 2.7]).
@@
For the case of $2g=2+n>0$, substitute each $x_{i}$ in (1.10) with $\begin{cases}x_{i}=z_{i}+\frac{1}{z_{i}},\\ z_{i}=\frac{t_{i}+1}{t_{i}-1},\end{cases}$ and write the result as $F_{g,n}(t_{1},\dots,t_{n})$ incorporating these substitutions. Then for every $(g,n)$ in the range of 2g2+n>02g-2+n>0, the following holds:
  • $F_{g,n}(t_{1},\dots,t_{n})$ is a Laurent polynomial in the $t$-variables of degree 6g6+3n6g-6+3n.
@@
Here, ${\mathcal{M}}_{g,n}$ is the moduli space of smooth $n$-pointed algebraic curves of genus $g$, ${\overline{\mathcal{M}}}_{g,n}$ its compactification, i.e., the Deligne-Mumford stack of stable curves of finite type $(g,n)$, and
@@
Although we do not touch in these lectures, there is yet another coordinate change from $(t_{1},t_{2},t_{3},\dots)$ appearing in the Catalan case to the time evolution parameters $(t_{1},t_{2},t_{3},\dots)$ in the Lax equations (1.17) defined below. Our use of the same notations is due to the (confusing) conventions of different previous publications. In reality, these are not the same variables. After the choice of the right coordinate transformation,
@@ 1.2. Theme $2$: When is a space a moduli space?
A moduli space is different. On this space, every point has its unique name. Each point is distinguished by its individuality. No points on a moduli space are the same. A neighborhood of each point is different from place to place. There may also be singularities on a moduli space. Each point has its own characteristic importance, yet it is a member of a family. Since every point is different, the moduli space itself presents its miracle: it can exist, as a harmonious space!
@@
We give a linear coordinate system on ${\mathbb{C}}^{n}$ and represent the $2n$ generators of $\Gamma$ by a set of $2n$ column vectors $w_{1},\dots,w_{2n}$ of size $n$. They form a complex matrix $\Pi$ of size $n\times 2n$. $GL_{n}({\mathbb{C}})$ acts on $\Pi$ from the left, representing the coordinate change. Since the column vectors of the matrix $\Pi$ generate the $\mathbb{Z}$}-module $\Gamma$, $SL_{2n}({\mathbb{Z}})$ acts on $\Pi$ from the right, representing change of generators of the same lattice $\Gamma$. Since $\{w_{1},\dots,w_{2n}\}$ is linearly independent over $\mathbb{R}$}, there is a $\mathbb{C}$}-linearly independent subset consisting of $n$ vectors. Thus we can use a left multiplication of $GL_{n}({\mathbb{C}})$ and a right multiplication of $SL_{2n}({\mathbb{Z}})$ on $\Pi$ to bring it to the special shape $I\,|\,\Omega]$. The geometry of the quotient space ${\mathbb{C}}^{n}/\Gamma$ is encoded in this $n\times n$ matrix $\Omega$, which we call the period matrix of the torus $T={\mathbb{C}}^{n}/\Gamma$.
@@
is the moduli space of holomorphic line bundles on $C$ of degree 0. For each fixed $g>1$, there are 3g33g-3 dimensional family of Jacobian varieties, while Abelian varieties form a family of $g(g+1)/2$ dimensions. Therefore, Jacobians are very special among Abelian varieties.
@@ 1.3. Theme $3$: Finding a curve from a period matrix
(1.16) $\vartheta(z,\Omega)=\sum_{n\in{\mathbb{Z}}^{n}}\exp\left(2\pi i{\langle}n,z{\rangle}+\pi i{\langle}n,\Omega n{\rangle}\right),\hskip 23.49976ptz\in{\mathbb{C}}^{n},$
where $\langle$$\;\;{\rangle}$ is the standard symmetric form on ${\mathbb{C}}^{n}$. The Riemann period condition makes the above infinite series absolutely convergent on ${\mathbb{C}}^{n}$ and satisfy quasi-periodic conditions.
@@
(1.17) $\frac{\partial L}{\partial t_{n}}=[B_{n},L],\hskip 23.49976ptB_{n}=(L^{n})_{+},\qquad n=1,2,3,\dots,$
@@
(1.18) $L=\partial+\sum_{i=1}^{\infty}u_{i+1}(x;\mathbf{t})\partial^{-i},\hskip 23.49976pt\partial=\frac{\partial}{\partial x},\qquad\mathbf{t}=(t_{1},t_{2},t_{3},\dots)$
@@
which does not depend on $y$ gives a solution to the KP equation. A relation between the KdV equation (1.19) and algebraic curves is easily seen by considering a traveling wave solution with speed c-c constructed from the Weierstraß  elliptic function $\wp(z)$:
@@ 1.4. Theme $4$: The $\tau$-functions
on the Sato Grassmannian, where $V={\mathbb{C}}((z))$ and $\Lambda\in\mathfrak{gl}_{\infty}$ is the maximal nilpotent element corresponding to the multiplication of $z^{-1}$ on $V$ (see for example, [74, Chapter 7]). The $\tau$-function records this action through the vantage point of the determinant line bundle. This also explains why so many soliton equations are exactly solved in terms of determinant. Since the initial value problem of the KP equations is uniquely solvable ([72, 73]), the space of solutions to the KP equations is identified with $Gr(\mu=0)$.
@@
the space of meromorphic sections of $\mathcal{L}$} with arbitrary poles at $p$, where $p\in C$ is a non-singular point of $C$, $z$ is a local parameter of $C$ around $p$, and $\mathcal{L}$} is a line bundle on $C$ of degree g(C)1g(C)-1. We then have (see [72, Lemma 3.7])
@@
2. The spectral curve for Hurwitz numbers
@@ 2.1. The Laplace transform
This conversation took place in the lobby of MSRI, Berkeley, after the end of the day. We shook hands tightly and happily, as we started to depart. We then noticed that there was a young mathematician listening to our conversation. At the end, he shouted:
@@
A mathematical picture has emerged in the last two decades since the discoveries of Eynard-Orantin [38], Mariño [69], Bouchard-Klemm-Mariño-Pasquetti [14] and Bouchard-Mariño [15] in physics, and many mathematical efforts including [13, 18, 34, 35, 37, 41, 65, 76, 77, 79, 80, 83, 84, 85, 86, 87, 93]. As a working hypothesis, we phrase it in the form of a principle.
@@
There are many examples of mathematical problems that fall in to this principle. Among them is the theory of simple Hurwitz numbers [37, 80] that we are going to present now. Besides Hurwitz numbers, numerous mathematical results have been established. They include counting of Grothendieck’s dessins d’enfants [18, 76, 83, 86], higher genus Catalan numbers [32, 34], orbifold Hurwitz numbers [13], double Hurwitz numbers and higher spin structures [78], and the stationary Gromov-Witten invariants of ${\mathbb{P}}^{1}$ [35, 87]. One of the most important results in the Gromov-Witten setting is the remodeling conjecture [14] and its solution [41] mentioned above. A new direction was suggested in Kontsevich-Soibelman [59]. It is now impossible to list all research in this direction, simply called topological recursion, so we refer to recent papers by many authors and the papers they cite: Andersen, Borot, Bouchard, Chidambaram, Do, Dunin-Barkowski, Eynard, Garcia-Failde, Iwaki, Lewański, Norbury, Orantin, Osuga, Shadrin, … A comprehensive and the most recent introduction is given by Bouchard [12], in which one can see how the focus of the topological recursion community has changed into new developments in the last decade.
@@
2.2. Simple Hurwitz numbers
In this subsection, we define simple Hurwitz numbers, and a combinatorial equation that they satisfy, the cut-and-join equation [48, 94], is proved. In the next subsection, the Laplace transform of the Hurwitz numbers is presented. The Laplace transformed holomorphic functions live on the mirror $B$-side of the model, according to Princeple 2.3. The Lambert curve is defined as the domain of holomorphy of these holomorphic functions. Then in the following subsections, we give the Laplace transform of the cut-and-join equation. The result is a simple polynomial recursion formula and is equivalent to the Eynard-Orantin topological recursion for the Lambert curve. We also give a straightforward and simple derivation [80] of the Witten-Kontsevich theorem on the $\psi$-class intersection numbers [24, 57, 98], and the $\lambda_{g}$-formula of Faber and Pandharipande [39, 40], using the recursion formula we establish in [80].